For any two subsets A and B of a set U , we define their symmetric difference to
be A ∆ B = (A \ B) ∪ (B \ A)
i) Check whether ∆ distributes over ∩ .
ii) Show that A∆ fy =A
iii) Prove that A ∆ B =(A ∩ B compliment ) ∪(A compliment ∩ B).
1
Expert's answer
2020-05-24T20:07:07-0400

1) Ans= "\Delta" is not distribute over "\cap" .

Counterexample : Let "A=\{ 1,2,3,4,5 \}" ,"B=\{ 4,5 ,6,7,8\}" ,

"C=\{ 7,8,9,10\}"

"B\cap C=\{ 7,8\}" .

"A \Delta (B\cap C)=(A - (B\cap C )) \cup ((B\cap C)-A)"

"=\{1,2,3,4,5\} \cup \{ 7,8 \}"

"=\{ 1,2,3,4,5,7,8\}" .

"A\Delta B=(A-B)\cup (B-A)=\{ 1,2,3\} \cup \{6,7,8\}"

"=\{ 1,2,3,6,7,8\}"

"A\Delta C=(A-C) \cup (C-A)=\{ 1,2,3,4,5\} \cup \{ 7,8,9,10\}"

"=\{ 1,2,3,4,5,7,8,9,10\}"

"(A\Delta B) \cap (A\Delta C)= \{ 1,2,3,7,8\}"

"\therefore \ A\Delta (B\cap C) \neq (A\Delta B) \cap (A\Delta C)."


2) ans : "A \Delta \phi=( A-\phi ) \cup (\phi - A)" "=A\cup \phi=A"

3) ans:= Claim : "A\Delta B=(A\cap B') \cup (A'\cap B)"

Let "x\in A-B \iff x\in A \ and \ x\notin B \iff x\in A \ and \ x\in B'"

"\iff x\in A\cap B'" .

Thus , "A-B=A\cap B'"

Similarly , "B-A= B\cap A'" .

Hence , "A\Delta B=(A\cap B') \cup (B\cap A')" .